when dividing exponents with the bases of negative powers what do i do?
example-
5^-1/10

Answer by chessace(471)

In intermediate calculations, you often don't have to do anything, just keep doing the appropriate amrithmetic on the exponents.
At the end, if you need to clean it up, actually compute it, etc:
1. Get rid of the "-" exponent by reciprical:
5^-1/10 = 1 / (5^1/10)
2. Get rif of the exponent denominator by taking a root:
5^1/10 = 10th root of 5 (might have to use logs)
Merge: ans = 1 / 10thRoot(5).